OpenStudy (anonymous):

Solve for the d L+ a + 7d Solve fot t l = prt Solve for c p= a+b+c

OpenStudy (skullpatrol):

The addition, subtraction, multiplication and division properties of equality guarantee that a. adding the same real number to or subtracting the same real number from equal numbers gives equal results b. multiplying or dividing equal numbers by the same nonzero real number gives equal results

OpenStudy (anonymous):

L = a + 7d L - a = 7d (L - a)/7 = d I = prt I/pr = t p = a + b + c p - a - b = c p - (a+b) = c